OpenStudy (anonymous):

Differentiate F(x) = (log9 x)7

OpenStudy (kc_kennylau):

You mean \(\log_9x\) or \(\log9x\)?

OpenStudy (anonymous):

the first one

OpenStudy (kc_kennylau):

Some log formula: \(\Large\frac{\log_na}{\log_nb}=\log_ba\)

OpenStudy (kc_kennylau):

In this case you'd like to transform it to \(\Large\left(\dfrac{\ln x}{\ln 9}\right)^7\)

OpenStudy (kc_kennylau):

Chain rule?

OpenStudy (anonymous):

so the final answer would be (ln x/ln 9)^7?

OpenStudy (kc_kennylau):

No, this is only the question

OpenStudy (anonymous):

I'm kinda lost right now but ok

OpenStudy (kc_kennylau):

\[\Large(\log_9x)^7=\Large\left(\dfrac{\ln x}{\ln 9}\right)^7\]

OpenStudy (anonymous):

oh ok. i see now. thanks

OpenStudy (kc_kennylau):

Continue using chain rule?

OpenStudy (wolf1728):

Benji - did you check my last answer (half-life)?

OpenStudy (anonymous):

yes wolfman1728. i figured how to do it. thanks for the help you guys

OpenStudy (wolf1728):

Okay - and sorry for the interruption.

OpenStudy (kc_kennylau):

Treat \(\dfrac{\ln x}{\ln9}\) as one thing first, now differentiate it